OpenStudy (anonymous):

Find the remainder when P(x)= x^4 -3x^3 +4x^2 +5 is divided by x+2

OpenStudy (anonymous):

when I did it I got 61, but if others would like to confirm this..long time since I done long division.

OpenStudy (anonymous):

\[(x+2)(x ^{3} -5x ^{2}+14x-28)\]

OpenStudy (anonymous):

Thank you! You're probably right. Would this work with synthetic as well?

OpenStudy (anonymous):

Dont understand the question..explain?

OpenStudy (anonymous):

Would you be able to divide using synthetic division?

OpenStudy (anonymous):

I dont know what synthetic division means...is it like long division...I've been doin maths for some time now, and I've never heard of this term.

OpenStudy (anonymous):

Nevermind then. It's Algebra II. D:

OpenStudy (anonymous):

I have no problem with algebra, give me a question and I will tell you how to do it.

OpenStudy (anonymous):

and u are a noob^

OpenStudy (anonymous):

and u are a noob^

OpenStudy (anonymous):

is that question for me joeymason or for ginamercedes

OpenStudy (anonymous):

Do you know how to do this one? Find the equation of the perpendicular bisector of the line connecting P(-4, 6) and Q(4, -2).

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the equation of a line: \[y _{2}-y _{1}=m(x _{2}-x _{1})\]

OpenStudy (anonymous):

y2=-2, y1=6, x2=4, x1=-4

OpenStudy (anonymous):

now for m(the slope) \[m=y _{2}-y _{1}/x _{2}-x _{1}\]

OpenStudy (anonymous):

-2-6 / 4+4 So -8/8 -1?

OpenStudy (anonymous):

if you fill in the values in these equations you will have the line that connects your two points. the perpendicular bisector will be a line which is perpendicular. you already know the slope of your line..lets say b/a, then the slope of the perpendicular line is -a/b

OpenStudy (anonymous):

So from here, I just plug it in?

OpenStudy (anonymous):

now, the perpendicular line is the opposite slope

OpenStudy (anonymous):

so if your slope is a/b, then the slope of the perpendicular line is -b/a

OpenStudy (anonymous):

sorry for the delay my system crashed had to reboot

OpenStudy (anonymous):

It's okay, thank you for your assistance thus far.

OpenStudy (anonymous):

now you have the slope of the perpendicular bisector, you need to find a point on that line

OpenStudy (anonymous):

to do this we find the midway point (hence the term bisector) which means cut in half

OpenStudy (anonymous):

the midway point is simply \[x _{1}+x _{2}/2, y _{1}+y _{2}/2\]

OpenStudy (anonymous):

once you have that you have the slope and one point on the line, you can then get the equation of the line

OpenStudy (anonymous):

in this case as you said the slope is -1

OpenStudy (anonymous):

or -1/1 therefore the slope of the perpendicular is -(1/-1) = 1

OpenStudy (anonymous):

now the bisector (point) is -4+4/2=0/2=0 and 6+(-2)=4 so your midpoint (the point on your line is (0,4)

OpenStudy (anonymous):

then putting this into the equation y-y1=m(x-x1) we get the following

OpenStudy (anonymous):

y-4=1(x-0) or y-4=x which becomes: y=x+4

OpenStudy (anonymous):

I hope that answers your question!

OpenStudy (anonymous):

Thank you so much!!!